OpenStudy (anonymous):

Ben has a 40% chance of being selected as the president of the school sports club and a 70% chance of being elected editor of the school magazine. What is the probability that Ben will be selected as the president of the sports club and will also be elected editor of the school magazine? A) 0.3% B) 1.1% C) 14% D)28%

OpenStudy (anonymous):

70 / 40 = 1.75 (not even close) So it's 70/ 40 % = 28 It's D When you calculated this, did you add a percent %?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh, well next time just make sure add the percent :)

OpenStudy (amistre64):

how does that process get us to 28% ?

OpenStudy (anonymous):

so you would take the 70 and divide it by 40

OpenStudy (anonymous):

wait no I still don't get it

OpenStudy (anonymous):

70 / 40 % = 28. Try it with your calculator :)

OpenStudy (amistre64):

theres a rule: P(A n B) = P(A) * P(B) if A and B are independant

OpenStudy (anonymous):

oh 40% so I would divide 70 by 0.4

OpenStudy (amistre64):

Maliboo is not making any sense to me ....

OpenStudy (anonymous):

still don't get it

OpenStudy (amistre64):

heres the rule, again: P(A n B) = P(A) * P(B) if A and B are independant good luck with it

OpenStudy (anonymous):

Guys, if you have a calculator please do 70 divided by 40 %, BUT YOU MUST MAKE SURE YOU ADD THE PERCENT AFTER 40 or it won't be 28

OpenStudy (anonymous):

so would 70 and 40 be A n B

OpenStudy (amistre64):

.70 and .40 yes

OpenStudy (anonymous):

When I did 70/40% I got 175

OpenStudy (amistre64):

dont follow the advice of a fool ...

OpenStudy (anonymous):

What is P then

OpenStudy (amistre64):

P(A) means: the probability of event A

OpenStudy (amistre64):

what class are you working this in?

OpenStudy (anonymous):

math

OpenStudy (amistre64):

math is a broad subject. you should have been exposed to some of this notation during the class at least

OpenStudy (amistre64):

define our events, what are we looking for?

OpenStudy (anonymous):

if I do 0.40 * 0.70 I get 0.28 so @Maliboo was correct

OpenStudy (amistre64):

her answer was right, her process was garbage

OpenStudy (anonymous):

hey thats not very nice to say there could be multiple methods

OpenStudy (anonymous):

Well, when I did it on MY CALCULATOR I got 28.

OpenStudy (amistre64):

no, there arent multiple ways of processing this. 70/40 is not .28 no matter what buttons you push on a calculator

OpenStudy (anonymous):

I unfanned you amistre64

OpenStudy (amistre64):

the next type of question you get like this .. you want to work the correct process, not take some strangers willy nilly nonsense that they cant even perform.

OpenStudy (amistre64):

good luck then, its your grade ....

OpenStudy (anonymous):

THANK YOU @Maliboo

OpenStudy (anonymous):

Your welcome :)

OpenStudy (amistre64):

spose someone process was: i took the retrograde of mars with jupiters orbital apex, and got .28 gets .28 but the rest is not in any way a reasonable assessment of the information. just saying :)